OpenStudy (anonymous):

Hello,I have done the 1A Exercise and I got a problem in the 1a-6-a)Expressing(sinx+√3cosx) in the form Asin(x+c),the solution given was 2sin(x+π/3),but I thought it could be 2sin(x-2π/3),Is it could be 2sin(x-2π/3)??

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

i expect you used the compound angle formula to do this problem and got Asin (x + c) = Asin x cos c + A cos x sinc and compare this identity to A sin x + A cos x * √3 so A = √(1 +3) = 2 and tan c = √3 so c = π/3 giving 2 sin (x + π/3)

OpenStudy (anonymous):

I got your point,but if I use the method like following I got the different answer: As sinx+√3cosx=asinkx+bcoskx,then a=1,b=√3,k=1,A=√(a^2+b^2)=√(1+3)=2,tankO(fai)=-b/a=-√3,then sinx+√3cosx=2sin(x-2/3π)

OpenStudy (dumbcow):

i agree with jimmyrep. however tan(pi/3) = tan(-2pi/3) therefore your answer is technically correct as well

OpenStudy (dumbcow):

wait nevermind, A would have to be -2 though

OpenStudy (dumbcow):

im not following your method...

OpenStudy (anonymous):

this is not a different answer this is the same answer